OpenStudy (anonymous):

Simplify the expression 1/3n(-6+27m-51p)

OpenStudy (anonymous):

i got = -2n + 9mn - 17pn showing my work 1/3n(-6+27m-51p) = -6/3n +27/3 mn -51/3 pn = -2n + 9mn - 17pn is it right @shubhamsrg

OpenStudy (shubhamsrg):

yesh! absolutely correct! ^_^

OpenStudy (anonymous):

thanks